OpenStudy (bmxlifestyle):

find the coordinate of the circumcenter ABC with vertices A (0,0) B (3,0) C (3,2)

OpenStudy (bmxlifestyle):

@Will.H

OpenStudy (bmxlifestyle):

A (0,0) B (1.5,1) C (2.5,1.5) D (6,2)

OpenStudy (bmxlifestyle):

@563blackghost can u help

OpenStudy (holsteremission):

The circumcenter is the center of the circle passing through the given vertices. Consider the general equation for a circle centered at \((a,b)\) with radius \(r\): \[(x-a)^2+(y-b)^2=r^2\]If you plug in each vertex's coordinate for \((x,y)\) in the above equation, you end up with the system \[\begin{cases} a^2+b^2=r^2&\text{when }(x,y)=(0,0)\\[1ex] (a-3)^2+b^2=r^2&\text{when }(x,y)=(3,0)\\[1ex] (a-3)^2+(b-2)^2=r^2&\text{when }(x,y)=(3,2) \end{cases}\]which can be solved for \(a,b,r\), but you only need the first two to find the center.

OpenStudy (bmxlifestyle):

so would the answer be A (0,0)

OpenStudy (holsteremission):

Have you tried solving the system? I'm not here to hand out answers.